if 2.61g of CuNO3 is dissolved in water to make a 0.480M solution, what is the volume of the solution?
1
Expert's answer
2019-09-23T03:47:44-0400

Moles of "CuNO_3=\frac{mass}{molar\ mass}=\frac{2.61}{125.55}=0.02078\ mole"


Molarity = "\frac{moles}{volume(in\ l)}=0.480=\frac{0.02078}{volume(in\ l)}"


volume = "\frac{0.02078}{0.480}=0.0433\ l"


volume"=43.3\ ml"